# Machine Learning for Structural Biology

### v1

@@ B.1 Substitution Matrix-based Variant Generation
By extracting the CDR regions from our initial dataset, we established the foundational alignments for our statistical analysis and subsequent construction of the substitution matrix, following the approach of Henikoff et al. [17]. In particular, we calculated substitution frequencies across homologous CDR regions using the Needleman-Wunsch alignment algorithm [28] to create a custom matrix that emphasizes biologically relevant mutations. From our curated CDR alignments, we counted every observed amino acid substitution background frequencies, and normalized the results to create a probabilistic framework for residue changes. The described stochastic matrix reflects the true mutational landscape of antibody hypervariable loops and defines the probability distribution that guides our sequence updates.
@@ B.2 Bioinformatics
The primary aim of Bioinformatics is to narrow the vast sequence space by filtering out the sequences that do not require further analysis. Since all the proceeding algorithm is initiated from the same analysis, we started by exploring and cleaning the sequence dataset received from Avery (Jewett et al. (2023) [20]). We used Riot-NA [12] to annotate full-length VH sequences and precisely isolate the three complementarity-determining regions (CDR-H1, H2, H3). By focusing solely on these regions, we addressed the key challenge of reducing the sequence space while retaining the most critical information from the antibody: the areas where binding with the target occurs. Riot-NA matches each sequence to a germline database to find the exact CDR regions. Then, we use strict filters to keep only the clear, high-quality CDRs for further analysis.
After a new mutation is produced, it goes through two steps of verification. Firstly, we verify that the new sequence is biologically feasible and can be produced experimentally. Inspired by [9] and [32], we introduced biophysical developability constraints to ensure that each mutated sequence is both synthetically accessible and capable of folding and binding effectively. Our algorithm enforces four principal criteria for developability: hydrophobicity (to promote the burial of hydrophobic residues within the antibody core), aromatic content (to enhance CDR stability and binding through aromatic interactions), charge (to avoid extreme net charges that can lead to aggregation or misfolding), and sequence patterns (to eliminate undesirable motifs such as N‑glycosylation sites or stretches of more than five identical residues).
The second step in our evaluation is structural verification using AlphaFold2, a state-of-the-art deep learning model that predicts full atomic protein structures from primary sequences [23]. Each candidate sequence is processed to generate high-confidence 3D models along with per-residue confidence metrics (pLDDT, pTM, ipTM). As the project evolved, AlphaFold2 was replaced with ImmuneBuilder, which, while lacking explicit cutoff metrics, employs an ensemble-based prediction strategy that effectively mitigates structural variability and improves robustness in downstream analysis.
@@ Appendix C Data Generation
We build a fully autonomous, end‑to‑end in silico pipeline by seamlessly integrating multiple independent, open‑source tools—Riot‑NA, ImmuneBuilder, PBDFixer, and HADDOCK3. Our Scoring Workflow start with each antibody sequence being first analyzed by Riot‑NA to delineate the heavy and light chain CDR loops. The trimmed sequences are then fed into ImmuneBuilder, which outputs both per‑residue confidence scores and full atomic 3D models. Any gaps or missing atoms in these models are automatically repaired by PBDFixer, ensuring a complete structure for downstream analysis. Finally, the corrected structures are sent to HADDOCK3, where physics‑based docking simulations produce quantitative docking scores that inform the selection of the most promising variants.
@@
We use the SARS‑CoV‑2 antibody database compiled by Jewett et al.[20], which catalogs experimentally characterized variable‑region sequences. Notably, the raw dataset provides improperly annotated CDR regions, which we re-annotate using Riot-NA [12]. Accurate identification of CDRs is essential for evaluating how sequence changes may affect protein–protein interactions, which is the foundation of our mutation and scoring workflow.
@@
In earlier versions of the pipeline, each candidate sequence was processed with AlphaFold2 [23] to generate predicted protein structures and associated confidence metrics (pLDDT, pTM, ipTM). Variants not meeting predefined thresholds for these scores were discarded prior to docking. This step was later replaced by ImmuneBuilder [1], which predicts antibody structures with state-of-the-art accuracy using deep learning models trained specifically on immune system proteins. The output is a set of four different predicted conformations, each in PDB format, which specifies the three-dimensional coordinates of each amino acid residue. Additionally, ImmuneBuilder outputs root mean square deviation (RMSD) values for each amino acid, which, intuitively, quantifies how far a given predicted amino acid is from the average predicted position of that amino acid across the four outputted structures.
@@
We perform the docking step using HADDOCK3 [19], a data-driven platform for modeling biomolecular complexes. HADDOCK3 integrates experimental data and physicochemical parameters to generate structural models of protein-to-protein interactions. For each antibody-antigen pair, the software produces multiple docked complexes in PDB format, along with detailed scoring information that includes van der Waals, electrostatics, desolvation energy, and an overall HADDOCK binding score. The final HADDOCK binding score is a weighted sum of energy terms, designed to approximate the quality and stability of the predicted protein–protein complex. A more negative score indicates stronger binding for a given complex, which is why we employ a minimization function in out set up. HADDOCK3 groups resulting models into clusters based on structural similarity, and only top-scoring clusters are retained for downstream analysis. This procedure enables us to prioritize variants with strong predicted binding ability and plausible interaction interfaces.
Appendix D Experimental
D.1 Hyperparameter Selection in Directed Evolution and Active Learning Approaches
@@
D.2 More on Ensemble Docking
@@ E.5 Results
Figure 5: Sidechain geneology and the resulting parent chain.

### v2

@@ B.1 Substitution Matrix-based Variant Generation
By extracting the CDR regions from our initial dataset, we established the foundational alignments for our statistical analysis and subsequent construction of the substitution matrix, following the approach of Henikoff et al. [henikoff1993performance]. In particular, we calculated substitution frequencies across homologous CDR regions using the Needleman-Wunsch alignment algorithm [murata199022] to create a custom matrix that emphasizes biologically relevant mutations. From our curated CDR alignments, we counted every observed amino acid substitution background frequencies, and normalized the results to create a probabilistic framework for residue changes. The described stochastic matrix reflects the true mutational landscape of antibody hypervariable loops and defines the probability distribution that guides our sequence updates.
@@ B.2 Bioinformatics
The primary aim of Bioinformatics is to narrow the vast sequence space by filtering out the sequences that do not require further analysis. Since all the proceeding algorithm is initiated from the same analysis, we started by exploring and cleaning the sequence dataset received from Avery (Jewett et al. (2023) [hunt2023rapid]). We used Riot-NA [riot] to annotate full-length VH sequences and precisely isolate the three complementarity-determining regions (CDR-H1, H2, H3). By focusing solely on these regions, we addressed the key challenge of reducing the sequence space while retaining the most critical information from the antibody: the areas where binding with the target occurs. Riot-NA matches each sequence to a germline database to find the exact CDR regions. Then, we use strict filters to keep only the clear, high-quality CDRs for further analysis.
After a new mutation is produced, it goes through two steps of verification. Firstly, we verify that the new sequence is biologically feasible and can be produced experimentally. Inspired by [csi-greifflab_developability_profiling] and [sabpred_tap], we introduced biophysical developability constraints to ensure that each mutated sequence is both synthetically accessible and capable of folding and binding effectively. Our algorithm enforces four principal criteria for developability: hydrophobicity (to promote the burial of hydrophobic residues within the antibody core), aromatic content (to enhance CDR stability and binding through aromatic interactions), charge (to avoid extreme net charges that can lead to aggregation or misfolding), and sequence patterns (to eliminate undesirable motifs such as N‑glycosylation sites or stretches of more than five identical residues).
The second step in our evaluation is structural verification using AlphaFold2, a state-of-the-art deep learning model that predicts full atomic protein structures from primary sequences [AlphaFold2021]. Each candidate sequence is processed to generate high-confidence 3D models along with per-residue confidence metrics (pLDDT, pTM, ipTM). As the project evolved, AlphaFold2 was replaced with ImmuneBuilder, which, while lacking explicit cutoff metrics, employs an ensemble-based prediction strategy that effectively mitigates structural variability and improves robustness in downstream analysis.
@@ Appendix C Data Generation
We build a fully autonomous, end‑to‑end in silico pipeline by seamlessly integrating multiple independent, open‑source tools—Riot‑NA, ImmuneBuilder, PBDFixer, and HADDOCK3. Our Scoring Workflow starts with each antibody sequence being first analyzed by Riot‑NA to delineate the heavy and light chain CDR loops. The trimmed sequences are then fed into ImmuneBuilder, which outputs both per‑residue confidence scores and full atomic 3D models. Any gaps or missing atoms in these models are automatically repaired by PBDFixer, ensuring a complete structure for downstream analysis. Finally, the corrected structures are sent to HADDOCK3, where physics‑based docking simulations produce quantitative docking scores that inform the selection of the most promising variants.
@@
We use the SARS‑CoV‑2 antibody database compiled by Jewett et al.[hunt2023rapid], which catalogs experimentally characterized variable‑region sequences. Notably, the raw dataset provides improperly annotated CDR regions, which we re-annotate using Riot-NA [riot]. Accurate identification of CDRs is essential for evaluating how sequence changes may affect protein–protein interactions, which is the foundation of our mutation and scoring workflow.
@@
In earlier versions of the pipeline, each candidate sequence was processed with AlphaFold2 [AlphaFold2021] to generate predicted protein structures and associated confidence metrics (pLDDT, pTM, ipTM). Variants not meeting predefined thresholds for these scores were discarded prior to docking. This step was later replaced by ImmuneBuilder [abanades2023immunebuilder], which predicts antibody structures with state-of-the-art accuracy using deep learning models trained specifically on immune system proteins. The output is a set of four different predicted conformations, each in PDB format, which specifies the three-dimensional coordinates of each amino acid residue. Additionally, ImmuneBuilder outputs root mean square deviation (RMSD) values for each amino acid, which, intuitively, quantifies how far a given predicted amino acid is from the average predicted position of that amino acid across the four outputted structures.
@@
We perform the docking step using HADDOCK3 [honorato2024haddock2], a data-driven platform for modeling biomolecular complexes. HADDOCK3 integrates experimental data and physicochemical parameters to generate structural models of protein-to-protein interactions. For each antibody-antigen pair, the software produces multiple docked complexes in PDB format, along with detailed scoring information that includes van der Waals, electrostatics, desolvation energy, and an overall HADDOCK binding score. The final HADDOCK binding score is a weighted sum of energy terms, designed to approximate the quality and stability of the predicted protein–protein complex. A more negative score indicates stronger binding for a given complex, which is why we employ a minimization function in our set up. HADDOCK3 groups resulting models into clusters based on structural similarity, and only top-scoring clusters are retained for downstream analysis. This procedure enables us to prioritize variants with strong predicted binding ability and plausible interaction interfaces.
Appendix D Experimental Procedures
D.1 Methodology
D.2 Hyperparameter Selection in Directed Evolution and Active Learning Approaches
@@
D.3 More on Ensemble Docking
@@ E.5 Results
Figure 5: Sidechain genealogy and the resulting parent chain.